OpenStudy (anonymous):

Use substitution to solve the system of equations. $y + 1 = x$ $-2x + 3y = 2$ a) (5, 4) b) (-6, 3/2) c) (1/2, 15) d) (1/4, 1)

OpenStudy (anonymous):

y + 1 = x -2x + 3y = 2

OpenStudy (anonymous):

**